### v1

# Using Large Language Models for Automatic Hint Ranking and Generation

Abstract
The use of Large Language Models (LLMs) has increased significantly recently, with individuals frequently interacting with chatbots to receive answers to a wide range of questions. In an era where information is readily accessible, it is crucial to stimulate and preserve human cognitive abilities and maintain strong reasoning skills. This paper addresses such challenges by promoting the use of hints as an alternative or a supplement to direct answers. We first introduce a manually constructed hint dataset, WikiHint, which includes 5,000 hints created for 1,000 questions. We then finetune open-source LLMs such as LLaMA-3.1 for hint generation in answer-aware and answer-agnostic contexts. We assess the effectiveness of the hints with human participants who try to answer questions with and without the aid of hints. Additionally, we introduce a lightweight evaluation method, HintRank, to evaluate and rank hints in both answer-aware and answer-agnostic settings. Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves hint quality, and (c) encoder-based models perform better than decoder-based models in hint ranking.
1 Introduction
Figure 1: Pipeline of WikiHint dataset generation. The numbers in arrows indicate the counts of output questions.
In recent years, question answering (QA) systems have risen in importance, giving users the opportunity to ask arbitrary questions and gain answers to them Mavi et al. (2024); Karpukhin et al. (2020); Zhao et al. (2021); Abdel-Nabi et al. (2023). The rapid development of Large Language Models (LLMs) Gemini Team et al. (2023); OpenAI et al. (2023); Dubey et al. (2024) has without doubt contributed to this, as well as to many other Natural Language Problems (NLPs) Qin et al. (2024). While the benefits of using LLMs for current information access are clear [^1] While the current LLMs have still considerable weaknesses such as hallucinations, once can assume that, in the future, these problems will be mitigated to large extent, judging from the speed of the recent technology advancement., there are some worries in relation to their potential effect on human development. One such concern relates to the potential weakening of important cognitive skills of users like thinking, reasoning, and remembering due to the expected widespread use of automatic question answering technologies, in particular, ones backed by powerful AI solutions  Heersmink (2024). Users who will rely mainly on the solutions presented by AI’s, might also be discouraged to practice and improve their reasoning abilities Alfredo et al. (2024). For example, Darvishi et al. (2024) demonstrate that students are more likely to depend on AI assistance instead of learning from it. Moreover, Jošt et al. (2024) examine the impact of LLMs as an automated problem-solving technology on education and learning outcomes, demonstrating that such systems can negatively affect the development of learning skills. Furthermore, psychological studies confirm the importance of obtaining an answer independently, enhancing user self-confidence and encouraging further learning Bandura (2013). Letting users come up with the correct answers by themselves should then also contribute to the positive psychological effect, potentially increasing their self-confidence and motivation for learning Usher and Pajares (2006).
While there is no simple remedy to the aforementioned problem, we would like to promote an approach that involves humans in the answer finding process, by providing them with hints rather than direct answers. Hints are meant to serve as subtle clues to guide potential users towards a correct answer, without revealing the solution Hume et al. (1996). This should engage human’s cognitive abilities, potentially leading to forming new pathways in brains based on the received hints and already possessed knowledge. Automatically generating hints for user questions could be used as an alternative for those who prefer to find the answers themselves; much like long-distance walking is an alternative to using a car for people who wish to take additional effort for the benefit of staying healthy Panter et al. (2018).
In this paper, we propose the first dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint which has been manually constructed and designed for both hint generation as well as evaluation. We next explore the performance of various LLMs in generating hints across different scenarios, including vanilla and finetuned models. We also examine the quality of the generated hints using both answer-aware and answer-agnostic approaches. Finally, we assess the effectiveness of a novel evaluation method for hint ranking called HintRank and compare it with other automatic evaluation techniques. To sum up, we make the following contributions in this paper:
  • We release the first manually created dataset called WikiHint for HG task containing 5,000 hints and 1,000 questions.
  • We propose an automatic evaluation method for ranking hints called HintRank and compare with other evaluation methods.
  • We finetune and evaluate LLMs on WikiHint to assess the dataset quality and LLMs capabilities in hint generation and ranking.
  • We present several novel observations, including the findings of a positive correlation between hint convergence and helpfulness, an inverse correlation between their length and helpfulness, and the superiority of the answer-aware approach over the answer-agnostic approach. In general, our research can contribute to fostering research in explainable AI for education Khosravi et al. (2022).
2 Related Work
Automatic question answering (QA) Karpukhin et al. (2020); Zhao et al. (2021); Abdel-Nabi et al. (2023) and question generation (QG) Kurdi et al. (2020); Lu and Lu (2021); Zhang et al. (2021) have advanced quite much in the last years. These tasks have seen numerous different datasets Trischler et al. (2017); Rajpurkar et al. (2016); Zhang et al. (2020), and evaluation metrics Nema and Khapra (2018); Mavi et al. (2022) proposed. The research related to hint generation is however still scarce despite that hinting is a common mechanism used by humans for question answering, and that automatic hint generation could be regarded as the third missing task alongside the two established ones, QA and QG. The prior research focused mainly on generating hints for programming Price et al. (2019); Kochmar et al. (2022); Barnes and Stamper (2008); McBroom et al. (2021) and typically in the context of intelligent tutoring systems.
Automatic hint generation for factoid questions was first addressed by Jatowt et al. (2023). However, the authors neither released a dataset nor utilized LLMs, focusing instead on hints generated from selected Wikidata [^3] https://www.wikidata.org/ predicates. Moreover, their work only considered an answer-aware setting and did not explore the hint ranking task. Subsequently, Mozafari et al. (2024) released the first synthetic dataset for hint generation (HG) called TriviaHG, which was automatically generated using LLMs. However, the automatic generation increases the likelihood of false information within the dataset, particularly due to the hallucination phenomenon of LLMs. The authors also introduced the first automatic evaluation method, called Convergence, for assessing hint quality. However, this method requires substantial computational resources as it relies on LLMs for evaluation. Interested readers can also refer to the recent survey of Jangra et al. (2024) who discuss various types of hints and challenges associated with hint generation and evaluation.
To the best of our knowledge, WikiHint is the first manually curated dataset for the HG task, with questions and hints verified by humans. Our proposed automatic evaluation method is also the first for this task that does not rely on LLMs and is lightweight enough to be used locally.
3 WikiHint Dataset
The absence of high-quality, verified hint datasets poses a significant challenge given the demanding data requirements of LLMs for their effective training. In this section, we outline the process for constructing WikiHint dataset. Figure 1 provides an overview of the pipeline of the dataset generation process, which we explore in detail in the following sections.
3.1 Question Sampling Module
We incorporated AI-generated questions using ChatGPT OpenAI (2022) and questions from existing popular QA datasets such as SQuAD 2.0 Rajpurkar et al. (2018) and Natural Questions (NQ) Kwiatkowski et al. (2019). The following prompt was used for question generation:
Can you give me 10 questions where the answer is ANSWER? Please put them in a CSV file with answer=ANSWER and link=WIKIPEDIA_LINK where each question has an answer and link. Make sure to put the questions in quotation marks.
where WIKIPEDIA_LINK is the URL for the Wikipedia page corresponding to ANSWER. As mentioned above, we also selected questions from SQuAD 2.0 and NQ making sure that their answers had dedicated Wikipedia articles with sufficiently long content. Finally, we manually verified the correctness of all the selected questions discarding questions that are either too general or have incorrect answers. Table 6 in Appendix A compares the difficulty levels of questions based on their sources.
Figure 2: The HintRank method.
3.2 Hint Creation
The crowdsourcing platform Amazon Mechanical Turk [^4] https://www.mturk.com/ was then used to distribute the hint creation task among multiple workers. The instructions shown to crowdworkers asked them to create five hints for a question and its associated Wikipedia article. After generating the hints, the workers were asked to rank them on a scale from 1 to 5, with 1 being the most helpful in finding correct answers and 5 being the least helpful. Figure 9 in Appendix A displays the annotators’ interface for hint generation while Figures 10, 11, and 12 provide additional views of the summarized instructions, detailed instructions, and the provided examples, respectively, to further assist the crowdworkers. Each data submission was subsequently reviewed manually for quality and was either approved or rejected based on their assessment. Table 14 in Appendix A shows the detailed criteria used for the selection process. The most common reasons for rejection were hints that directly revealed the answers (answer leakage) and hints that were single words instead of complete sentences. Among the 2,788 submissions reviewed, 1,788 were rejected and 1,000 were accepted. We prepared several attributes to be included for each question, answer, and hint, as shown in Tables 7, 8, and 9 in Appendix A. We discuss some of them in Section 4. Finally, we divided the questions and hints into train and test subsets. The train subset includes 4,500 hints for 900 questions, while the test subset contains 500 hints for 100 questions. Tables 10, 11, 12, and 13 in Appendix A show few examples of hints taken from the WikiHint dataset.
Train Test
Number of hints 4,500 500
Number of questions 900 100
Avg. question length (words) 19.55 19.19
Avg. hint length (words) 17.77 18.32
Avg. #entities / question 1.2 1.44
Avg. #entities / hint 1.2 1.18
Table 1: Statistics of WikiHint dataset.
Rank 1 2 3 4 5
Average Length 16.99 17.67 18.02 18.14 18.3
Table 2: Average length of hints vs. their ranks.
Dataset Subset Relevance Readability Convergence Familiarity Length
Answer Leakage
Degree (Avg)
Answer Leakage
Degree (Max)
TriviaHG Entire 0.95 0.71 0.57 0.77 20.82 0.23 0.44
WikiHint Entire 0.98 0.72 0.73 0.75 17.82 0.24 0.49
TriviaHG Train 0.95 0.73 0.57 0.75 21.19 0.22 0.44
WikiHint Train 0.98 0.71 0.74 0.76 17.77 0.24 0.49
TriviaHG Test 0.95 0.73 0.6 0.77 20.97 0.23 0.44
WikiHint Test 0.98 0.83 0.72 0.73 18.32 0.24 0.47
Table 3: Quality comparison of WikiHint and TriviaHG. Relevance, convergence, familiarity, and answer leakage are measured on a scale from 0 to 1, while readability is rated on a scale from 0 to 2 (the lower, the more readable).
4 Evaluation Approaches
Mozafari et al. (2024) proposed several evaluation metrics for assessing the quality of hints, including Relevance, Readability, Convergence, and Familiarity, although the authors proposed automatic evaluation methods only for Convergence and Familiarity. Convergence is a measure of how effectively a hint can narrow down or eliminate potential answers to a given question. Familiarity measures the expected level of knowledge of information expressed in hints. To evaluate hints based on Convergence and Familiarity, we follow the method proposed by Mozafari et al. (2024). However, we employ two cores for convergence including LLaMA-3.1-8b and LLaMA-3.1-70b Dubey et al. (2024).
We extend the above evaluation scheme by incorporating automatic methods for evaluating hint’s Relevance and Readability. We also propose a new metric for evaluating the probability of answer leakage - a case when a hint directly reveals the answer in its content. We introduce a lightweight automatic evaluation method for assessing hint quality in a pairwise scenario. The way to compute those additional metrics is briefly described below.
Hints can be considered a form of an answer since they provide explanations of the question’s correct answer. Based on this, one can evaluate the Relevance of a hint to its question as an Answer Relevance task Es et al. (2024) - the task where the goal is to assess how pertinent the provided answer is to the target question. To compute the answer relevance metric, we employ DeepEval framework [^5] https://docs.confident-ai.com/ treating hint as a kind of answer. To evaluate Readability Liu and Lee (2023), we finetune a RoBERTa Liu et al. (2019) model as a classifier on the OneStopEnglish dataset Vajjala and Lučić (2018). The finetuned model categorizes sentences into three classes: Beginner (0), Intermediate (1), and Advanced (2), reflecting their level of reading difficulty [^6] The accuracy of the readability estimator model is 62.3%.
To calculate the Answer Leakage Degree, we measure the semantic similarity between each word of a hint and an answer using RoBERTa model.
4.1 HintRank Evaluation Approach
In addition to the above automatic evaluation approaches involving individual hints, we introduce a new lightweight evaluation method, HintRank, for evaluating and ranking hints using pairwise preferences. Building on the success of widely-used automatic evaluation metrics like BERTScore Zhang et al. (2020), BEM Bulian et al. (2022), MoverScore Zhao et al. (2019), and BLEURT Sellam et al. (2020), which leverage BERT Devlin et al. (2019) as the core evaluation module and demonstrate its effectiveness, we chose BERT as the foundation for the HintRank method. Our method determines the better hint within a pair of hints. Figure 2 illustrates the proposed method. In the HintRank method, we begin by concatenating a given question and its answer with two hints, labeled as $\text{Hint}_{1}$ and $\text{Hint}_{2}$, to create an input compatible with BERT model. Note that in the answer-agnostic scenarios, we avoid appending the answer to the evaluated hints. Such constructed input is then processed by BERT model, which produces one of two possible outputs: 0 or 1. An output of 0 means that $\text{Hint}_{2}$ is of higher quality than $\text{Hint}_{1}$, whereas an output of 1 suggests that $\text{Hint}_{1}$ is superior to $\text{Hint}_{2}$. As HintRank operates on pairwise preferences, it requires $\binom{n}{2}$ comparisons for a question with $n$ hints, with a runtime complexity of $\mathcal{O}(n^{2})$.
5 Experiments and Results
5.1 Data Analysis
The WikiHint dataset is split into a train set with 4,500 hints (900 questions) and a test set with 500 hints (100 questions). Table 1 provides the statistics of both train and test sets, while Figure 7 in Appendix A shows their distributions according to the question types, indicating that the distributions are well-matched.
Figure 3: Average convergence of the hints of WikiHint based on the hint ranks.
We next analyze the difficulty levels of questions in WikiHint. To evaluate the difficulty, we utilize the Reference-based Question Complexity method Gabburo et al. (2024). This method computes the difficulty of a question by assessing how many of its retrieved passages contain the correct answer and by measuring the relevance between the retrieved passages and the question. It then calculates the difficulty score for the question based on such computed features. In particular, we use the DPR method Karpukhin et al. (2020) as the retrieval technique, employing an English Wikipedia dump preprocessed by Karpukhin et al. (2020) as the evidence source, and consider the top 30 most relevant passages as the retrieved passages. Figure 8 in Appendix A illustrates the computed question difficulty of WikiHint for train and test subsets [^7] We classify questions with difficulty scores below 0.33 as easy, those above 0.66 as hard and the rest as medium.. The figure indicates that medium-hard questions are the most common as well as the train and test subsets have quite similar distributions in terms of question difficulty. Also, Table 6 in Appendix A highlights the difficulty levels of questions generated or extracted from various sources.
Table 2 reveals an interesting insight regarding the length of hints, which can be considered as one of indicators of helpfulness. The results suggest that high-quality hints tend to be shorter in length (measured by the number of words) than the lower quality hints. This finding indicates an inverse correlation between hint length and helpfulness, challenging the intuition that longer hints are more informative or specific, and therefore more useful. In contrast, shorter hints appear to be more concise and easier to follow, likely presenting more helpful information in the first place.
We also evaluate the hints in the entire WikiHint dataset (and separately, in its train and test subsets) using the relevance, readability, convergence, familiarity, length, and answer leakage degree. We then compare these values with the ones obtained for TriviaHG dataset Mozafari et al. (2024) - the only existing hint dataset. The comparison results are presented in Table 3. The results indicate that in terms of relevance, readability, answer leakage degree, and familiarity, the metrics are nearly same between the two datasets. However, WikiHint has better convergence values compared to TriviaHG. Additionally, the hints in WikiHint are shorter in length, as measured by word count, than TriviaHG. These results indicate that the hints in WikiHint are of higher quality.
Lastly, Figure 3 demonstrates the negative correlation between the convergence scores of hints and their helpfulness as represented by hint ranks assigned by crowdworkers. The plot suggests that the convergence scores can be considered a reliable metric for evaluating the helpfulness of hints and for hint ranking.
5.2 Human Evaluation
To manually evaluate hints, we recruited five independent evaluators, who were not involved in the dataset generation process, to answer the questions from the test subset of the WikiHint. The process was as follows: 1. Participants were asked to answer the question without using any hints. If they provided a correct answer, they proceeded to the next question. 2. If they could not answer the question correctly, they were asked to review the hints until they could find the correct answer. By providing the correct answer, the participants could move to the next question. 3. If the participants could not answer the question after reviewing all the hints, they were allowed to skip the question.
Figure 4: The results of human evaluation.
Model Config
Use
Answer
Relevance Readability
Convergence
(LLaMA 8b)
Convergence
(LLaMA 70b)
Familiarity Length
Answer Leakage
Degree (Avg)
Answer Leakage
Degree (Max)
GPT-4 Vanilla 0.91 1.0 0.14 0.48 0.84 26.36 0.23 0.51
GPT-4 Vanilla 0.92 1.1 0.12 0.47 0.81 26.93 0.24 0.52
LLaMA-3.1-405b Vanilla 0.94 1.49 0.11 0.47 0.76 41.81 0.23 0.5
LLaMA-3.1-405b Vanilla 0.92 1.53 0.1 0.45 0.78 50.91 0.23 0.5
LLaMA-3.1-70b FTwA 0.88 1.5 0.09 0.42 0.84 43.69 0.22 0.48
LLaMA-3.1-70b Vanilla 0.86 1.53 0.05 0.42 0.8 45.51 0.23 0.5
LLaMA-3.1-70b FTwoA 0.86 1.5 0.08 0.38 0.8 51.07 0.22 0.51
LLaMA-3.1-70b Vanilla 0.87 1.56 0.06 0.38 0.76 53.24 0.22 0.5
LLaMA-3.1-8b FTwA 0.78 1.63 0.05 0.37 0.79 50.33 0.22 0.52
LLaMA-3.1-8b Vanilla 0.81 1.72 0.05 0.32 0.8 54.38 0.22 0.5
LLaMA-3.1-8b FTwoA 0.76 1.7 0.03 0.32 0.8 55.02 0.22 0.51
LLaMA-3.1-8b Vanilla 0.78 1.76 0.04 0.3 0.83 52.99 0.22 0.5
Table 4: Evaluation of generated hints based on relevance, readability, convergence, familiarity, length, and answer leakage across different scenarios.
Figure 4 illustrates that all the participants could answer more questions across all of question types such as HUMAN, ENTITY, and LOCATION [^8] We use names as stated in the original dataset. when they used hints compared to the case without hints. Notably, the greatest improvement was observed in human-related questions, where hints proved most beneficial. Following, entity-related questions led to significant improvement, while location-related questions saw the smallest positive change. This suggests that generating effective hints becomes progressively more challenging for human, entity, and location questions, in that order.
5.3 Model Performance
To further assess the quality of hints, we analyze how well LLMs can automatically generate hints for questions. We use the open-source LLaMA models: LLaMA-3.1-8b, LLaMA-3.1-70b, and LLaMA-3.1-405b Dubey et al. (2024), and GPT-4 OpenAI et al. (2023) as the most powerful closed-source LLM for comparison.
To explore different scenarios, we finetune [^9] We perform model finetuning using the API functions available on together.ai LLaMA-3.1-8b and LLaMA-3.1-70b on the training subset of the WikiHint dataset to evaluate the LLMs’ capabilities in hint generation when trained specifically on this task. For each question, we assign a hint as the target during the finetuning process. As a result of this learning strategy, during the inference stage, the finetuned model is prompted to generate one hint for each question. We consider two finetuning approaches: answer-aware and answer-agnostic. Given that LLMs typically handle most knowledge questions correctly, the answer-agnostic approach might be sufficient for generating hints. Besides, users generally do not know the answers to their questions when seeking hints. However, the answer-aware approach has its own advantages, too, such as in educational contexts where a teacher might use it to collect materials for class preparation. Due to the importance of both approaches, we chose to investigate fine-tuning of the LLMs in these two distinct scenarios.
We found that shorter prompts were more effective in achieving the desired task. Longer, more detailed instructions often led to the model disregarding the key goal, i.e., generating hints, and instead focusing on irrelevant details. In contrast, shorter prompts increased the likelihood of successful task completion. After experimenting, we opted for the following prompt as the system prompt:
You are a hint generator for the factoid questions. The user asks you a question and you should generate a hint for that question without revealing the answer in the hint.
Two distinct user prompts were employed to generate hints within a zero-shot learning strategy. Assuming a question $q$ as an input, the answer-agnostic prompt was ‘Give me the best hint for this question: $q$. The answer-aware prompt included the answer $a$ as follows: ‘Give me the best hint for this question: $q$? The answer for the question is $a$’.
Figure 5: Average length of generated hints by LLMs.
To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches: Vanilla-wA, Vanilla-woA, FTwA, and FTwoA where wA means With Answer and woA means Without Answer. We test these models on the WikiHint test to assess the impact of finetuning and the inclusion of answers in the prompt.
Figure 5 illustrates that as LLMs decrease in size and hint generation capability, the length of the generated hints increases. This supports our observation made in Section 5.1 of an inverse correlation between hint length and hint quality. Additionally, hints produced by finetuned models are generally shorter than those from vanilla models, indicating that finetuned models may generate higher-quality hints. Moreover, hints in the answer-aware scenarios are shorter compared to those in answer-agnostic scenarios, suggesting that when the answer is provided along with the question, LLMs are able to produce more effective hints.
Table 4 presents the quality of generated hints, evaluated with methods such as relevance, readability, convergence, answer leakage degree, and familiarity. The results indicate that more powerful LLMs are capable of generating more relevant hints. Regarding readability, GPT-4 exhibits the highest quality, followed closely by LLaMA-3.1-405b and LLaMA-3.1-70b, while LLaMA-3.1-8b shows the lowest readability. It also demonstrates that more powerful LLMs can generate more readable hints. Additionally, finetuned models consistently outperform their vanilla counterparts, and answer-aware prompts yield better results compared to answer-agnostic prompts for readability and familiarity. The answer leakage degree indicates that the prompt we use is effective in preventing LLMs from including answers, their synonyms or very similar terms in the generated hints, as the results closely align with those of WikiHint shown in Table 3.
Figure 6 illustrates that as LLMs decrease in size and capability, the convergence of their hints also diminishes. This trend is observed for both LLaMA-3.1-8b and LLaMA-3.1-70b, used as the cores of the convergence method. This supports our claim in Section 5.1 regarding the correlation between convergence and ranks. The figure also shows that the average convergence for the answer-aware approach surpasses that of the answer-agnostic approach, suggesting that including the answer in the prompt makes it easier for LLMs to generate hints. Furthermore, LLMs finetuned on the train subset of the WikiHint dataset achieve better convergence scores than their vanilla counterparts, indicating the efficacy of WikiHint for finetuning LLMs in hint generation.
Figure 6: Average convergence of the generated hints by different LLMs. The order of LLMs is determined by their capabilities and the parameter count.
5.4 HintRank Evaluation Method
Method Config Use Answer Accuracy (%) Correlation (%)
Convergence Vanilla 40.80 36.70
LLaMA-3.1-8b Vanilla 60.50 49.25
LLaMA-3.1-8b Vanilla 60.95 49.79
LLaMA-3.1-8b FTwoA 61.00 50.74
LLaMA-3.1-8b FTwA 61.25 49.03
LLaMA-3.1-70b Vanilla 64.00 50.32
LLaMA-3.1-70b Vanilla 64.25 51.32
LLaMA-3.1-70b FTwoA 64.65 51.51
LLaMA-3.1-70b FTwA 65.30 52.53
HintRank FTwoA 67.25 49.06
HintRank FTwA 68.55 52.34
Table 5: Comparison between Convergence metric, LLM-based ranking, and HintRank.
As outlined in Section 4, we also propose in this paper a novel evaluation method, HintRank, for ranking hints using the BERT model. Alongside finetuning BERT, we additionally finetune LLaMA-3.1-8b and LLaMA-3.1-70b models on the train set of the WikiHint to assess the performance of these LLMs in identifying high-quality hints. Similar to the experiments described in Section 5.3, we examine various scenarios including answer-aware and answer-agnostic contexts, and compare vanilla models with their finetuned counterparts. We use the following prompt as the system prompt:
You are a hint evaluator for the factoid questions. The user gives you a question and two hints and you should specify which hint for that question is a better hint and more helpful.
Two distinct user prompts are employed to evaluate hints within a zero-shot learning strategy. Assuming a question $q$ as a question and $h_{1}$ and $h_{2}$ as a pair of hints, the answer-agnostic prompt is:
Which hint is better to find the answer of this question: $q$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between "Hint_1" and "Hint_2" without any explanations.
and the answer-aware prompt with answer $a$ is:
Which hint is better to find the answer of this question: $q$. The answer for this question is $a$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between "Hint_1" and "Hint_2" without any explanations.
We benchmark HintRank against the Convergence metric which turned out to be useful for hint ranking assessment as indicated in Figure 3. To convert pairwise rankings to listwise rankings, we apply the Bradley–Terry model Bradley and Terry (1952). We evaluate the correlation between the rankings with Pearson Correlation Mining (2006).
Table 5 outlines the key features and differences among various scenarios. The results indicate that with the increase in the size and power of LLMs, both accuracy and correlation improve. Additionally, the answer-aware approach yields better outcomes compared to the answer-agnostic method, suggesting that the presence of an answer enables LLMs to evaluate hints more effectively. Moreover, finetuned versions outperform their vanilla counterparts, demonstrating that the WikiHint dataset is well-suited for model fine-tuning to rank hints. Surprisingly, the BERT-base method outperforms LLMs, including finetuned versions in the answer-aware scenario. This holds true for both Bert-FTwoA and Bert-FTwA, although BERT-base also performs better in the answer-aware approach compared to answer-agnostic. BERT-base methods achieve higher accuracy than LLMs and convergence, but in terms of correlation, LLaMA-3.1-70b exhibits the best performance. The effectiveness of BERT-base methods may be attributed to the strengths of encoder-based models like BERT in classification tasks over decoder-based models. Utilizing BERT-based models instead of LLMs enhances the speed and accessibility of HintRank, reducing computational demands. Figure 13 in Appendix A shows that accuracy improves as the rank difference between hints increases, indicating it’s harder to correctly order hints with closer ranks.
6 Conclusions
In this paper, we introduced the first manually created dataset for hint generation and hint ranking. We also presented a new lightweight method for evaluating and ranking hints. To demonstrate the effectiveness of our dataset, we conducted experiments where humans attempted to answer questions with and without the use of hints. The results confirm that the hints are of sufficient quality to assist users. We then finetuned LLMs using our dataset, prompting them to generate new hints for different questions. The high quality of the generated hints indicates that our dataset is well-suited for finetuning LLMs for HG task. We also finetuned BERT and LLMs on the dataset for the task of hint ranking and evaluated their performance. The results reveal that encoder-based models outperform decoder-based models in hint ranking.
In future, we plan to generate personalized hints tailored to the knowledge of askers. The main challenge here will be to develop appropriate datasets and solutions for user profiling.
Limitations
Our study has the following limitations:
  • The need for generative capabilities in hint generation task necessitates the use of LLMs. However, this dependency is a limitation as fine-tuning and prompting LLMs require extensive computational resources and are time-consuming.
  • Our research focus on factoid questions may limit its applicability to other types of questions that involve more complex or abstract answers. Factoid questions, by their nature, provide clear and concrete answers, which simplifies automated hint generation and evaluation but may not fully capture the breadth of human inquiry.
  • The WikiHint dataset is exclusively written in the English language. While this facilitates accessibility for a global audience and ensures compatibility with most existing Large Language Models, it also limits the dataset’s applicability in multilingual or non-English contexts, potentially excluding non-English speakers and diverse linguistic data.
Ethical Considerations
Our study utilizes GPT models, which are covered by the OpenAI License and Apache-2.0 license, and the LLaMA model, which is distributed under Meta’s LLaMA 2 Community License Agreement. We comply with these licensing agreements in all applications. Additionally, the datasets we use are sourced from repositories that are approved for academic use. The artifacts developed during our research are made available under the MIT license to facilitate straightforward modifications and use by the research community. We ensure that our data management, model training, and dissemination practices meet ethical standards and legal requirements associated with each artifact we use.
Appendix A Appendix
Figure 7: The distribution of Train and Test subsets.
Figure 8: Question difficulty based on different question types.
Source ChatGPT NQ SQuAD 2.0
Difficulty 0.43 0.34 0.38
Table 6: Question distributions based on the their sources and difficulty.
Figure 9: The MTurk Worker interface for the hint generation task.
Figure 10: The summarized instructions for the hint generation task.
Figure 11: The detailed instructions for the hint generation task.
Figure 12: Good and bad examples for the hint generation task.
Attribute Description
question The content of the question.
major The major category of the question.
minor The specific sub-category of the question.
entity Content of an entity
ent_type Type of entity (e.g., GPE, PERSON).
start_index Start index of the entity in the question.
end_index End index of the entity in the question.
wikipedia_page_title The title of the corresponding Wikipedia page to the entity in question.
wiki_views_per_month Number of views per month of the Wikipedia page for the entity in question.
normalized_views Views normalized to scale from 0 to 1 for the entity in the question.
readability Indicates the readability score
familiarity Indicates the familiarity score
difficulty Indicates the difficulty level of the question
Table 7: A detailed description of attributes of a question in WikiHint.
Attribute Description
answer The actual answer.
entity Content of an entity as identified within the answer.
ent_type Type of entity.
start_index Start index of the entity in the answer.
end_index End index of the entity in the answer.
wikipedia_page_title The title of the corresponding Wikipedia page to the entity in the answer.
wiki_views_per_month Number of views per month of the Wikipedia page for the entity in the answer.
normalized_views Views normalized to scale from 0 to 1 for the entity in the answer.
familiarity Indicates the familiarity score
difficulty Indicates the difficulty level of the answer
Table 8: A detailed description of attributes of an answer in WikiHint.
Attribute Description
hint Hint provided to assist with the question.
source URL source of the hint.
entity Content of entities mentioned in the hint.
ent_type Category of each entity (e.g., PERSON, GPE) mentioned in the hint.
start_index Specific start index where the entity is found in the hint text.
end_index Specific end index where the entity is found in the hint text.
wikipedia_page_title The title of the corresponding Wikipedia page to the entities in the hint.
wiki_views_per_month Number of views per month of the Wikipedia pages for the entities of the hint.
normalized_views Views normalized to scale from 0 to 1 for the entity in the hint.
relevance Indicates the relevance score
readability Indicates the readability score
convergence Indicates the convergence score
familiarity Indicates the familiarity score
answer_leakage Indicates the answer leakage score
rank Priority or helpfulness rating of the hint
Table 9: A detailed description of attributes of a hint in WikiHint.
Question What is the driest country in sub-Saharan Africa?
Fifth Best Hint The earliest settlers in this country were in the 18th century and crossed the Orange River to move into the area.
Fourth Best Hint The name of this country is derived from the oldest desert on Earth.
Third Best Hint This country’s coat of arms is a shield with the same design as the flag, 2 antelopes and a red blue and white bird.
Second Best Hint This country is a country in southern Africa with its western border along the Atlantic Ocean.
Best Hint This country’s flag is blue and green with a red stripe down the middle and the symbol of a sun in the top left corner.
Table 10: First example of the WikiHint.
Question What artist received the Presidential Medal of Freedom posthumously in 2018, in recognition of his contributions to American culture?
Fifth Best Hint The artist had karate as a lifelong interest, including its moves in his performances.
Fourth Best Hint The artist served in the Army after he became famous.
Third Best Hint The artist never performed outside North America.
Second Best Hint The artist tried to establish a career in films with pictures such as Jailhouse Rock and Fun in Acapulco.
Best Hint The artist is known for hits such as "Love Me Tender" and "Jailhouse Rock".
Table 11: Second example of the WikiHint.
Question What British comedy team is famous for its ’Four Yorkshiremen’ sketch, portraying exaggerated tales of hardship with humor?
Fifth Best Hint A giant cupid’s foot is repeatedly used in the show.
Fourth Best Hint This British comedy troupe is formed in 1969 consisting of 6 members.
Third Best Hint The team was awarded the AFI Star Award by the American Film Institute in 1998.
Second Best Hint The team received the BAFTA Award for Outstanding British Contribution to Cinema at the 41st British Academy Film Awards in 1988.
Best Hint The Holy Grail and Life of Brian are some of their greatest comedy films.
Table 12: Third example of the WikiHint.
Question What fictional pirate is known for carrying a compass that doesn’t point north but rather to what the user wants most?
Fifth Best Hint One of the pirate’s most famous quotes is "Now, bring me that horizon".
Fourth Best Hint The fictional pirate appeared in the video game series "Kingdom Hearts".
Third Best Hint The fictional pirate has a blood debt to another character called Davy Jones.
Second Best Hint The ship of the fictional pirate is called the Black Pearl.
Best Hint The fictional pirate is played by Johnny Depp in the film series.
Table 13: Fourth example of the WikiHint.
A hint must not include the exact answer explicitly. A hint must be a sentence. A hint must be specific, not generic. A hint must be from the corresponding Wikipedia page. A hint must have a unique rank.
Table 14: A detailed criteria or standards used for verifying the generated hints during the selection process.
Figure 13: Accuracy of HintRank for different hint pairs in different scenarios. The element at position $(r,c)$ represents the accuracy when comparing $\text{Hint}_{1}$ at rank $r$ to $\text{Hint}_{2}$ at rank $c$.